The actuator (pendulum) gains for the Beam Splitter and the two ITMs were measured to be:
BS: 9.54 +/- 0.05 nm/ct [100 ohm series resistor in coil driver board]
ITMX: 2.44 +/- 0.01 nm/ct [400 ohm series resistor in coil driver board]
ITMY: 2.44 +/- 0.02 nm/ct [400 ohm series resistor in coil driver board]
Counts here refers to DAC counts at the output of the coil filter banks (as opposed to counts at the LSC servo output). The dominant (systematic) uncertainty (which isn't quoted here) in this measurement is the determination of the peak-to-peak swing of the dark port sensor, AS55_Q. I estimate this error to be ~1ct/33cts = 3%. These values are surprisingly consistent with one another once we take into account the series resistance.
The last time this was done, we used ASDC to do the measurement. But I don't know what signal conditioning ASDC undergoes (in PD or in readout electronics). In any case, in my early trials yesterday, ASDC was behaving unpredictably. So I decided to do redo the measurement.
[Attachment #1]- Flowchart describing the calibration procedure.
[Attachment #2] - AS55_Q output while the Michelson was freeswinging. I had first aligned the ITMs using ASS. The peak-to-peak value of this corresponds to . So we know AS55_Q in terms of cts/m of MICH displacement.
[Attachment #3] - Magnitudes of transfer function from moving one of the MICH optics, to the now calibrated AS55_Q. Fits are to a shape , with being the fitted parameter. Coherence during the measurement is also plotted.
 - http://www.phys.ufl.edu/~bernard/papers/CQG20_S903.pdf